# ConceptAttention: Diffusion Transformers Learn Highly Interpretable Features

### v1

@@ Abstract
Do the rich representations of multi-modal diffusion transformers (DiTs) exhibit unique properties that enhance their interpretability? We introduce ConceptAttention, a novel method that leverages the expressive power of DiT attention layers to generate high-quality saliency maps that precisely locate textual concepts within images [^4] Code: https://alechelbling.com/ConceptAttention/. Without requiring additional training, ConceptAttention repurposes the parameters of DiT attention layers to produce highly contextualized concept embeddings, contributing the major discovery that performing linear projections in the output space of DiT attention layers yields significantly sharper saliency maps compared to commonly used cross-attention mechanisms. Remarkably, ConceptAttention even achieves state-of-the-art performance on zero-shot image segmentation benchmarks, outperforming 11 other zero-shot interpretability methods on the ImageNet-Segmentation dataset and on a single-class subset of PascalVOC. Our work contributes the first evidence that the representations of multi-modal DiT models like Flux are highly transferable to vision tasks like segmentation, even outperforming multi-modal foundation models like CLIP.
@@ 1 Introduction
Diffusion models have recently gained widespread popularity, emerging as the state-of-the-art approach for a variety of generative tasks, particularly text-to-image synthesis (Rombach et al., 2022). These models transform random noise into photorealistic images guided by textual descriptions, achieving unprecedented fidelity and detail. Despite the impressive generative capabilities of diffusion models, our understanding of their internal mechanisms remains limited. Diffusion models operate as black boxes, where the relationships between input prompts and generated outputs are visible, but the decision-making processes that connect them are hidden from human understanding.
@@
  • The novel discovery that the output vectors of attention operations produce higher-quality saliency maps than cross attentions. ConceptAttention repurposes the parameters of DiT attention layers to produce a set of rich textual embeddings corresponding to different concepts, something that is uniquely enabled by multi-modal DiT architectures. By performing linear projections between these concept embeddings and image patch representations in the attention output space we can produce high quality saliency maps.
  • ConceptAttention generalizes to achieve state-of-the-art performance in zero-shot segmentation on benchmarks like ImageNet Segmentation and Pascal VOC. We achieve superior performance to a diverse set of zero-shot interpretability methods based on various foundation models like CLIP, DINO, and UNet-based diffusion models; this highlights the potential for the representations of DiTs to transfer to important downstream vision tasks like segmentation.
  • We make ConceptAttention available, allowing researchers and practitioners to interpret and explore the intricate dynamics of text-to-image diffusion transformers. See code at: https://github.com/helblazer811/ConceptAttention.
@@ 2 Related Work
A variety of works train a diffusion models from scratch for the task of image segmentation (Amit et al., 2022; Karazija et al., 2024) or attempt to fine-tune pretrained models (Baranchuk et al., 2022). Another line of work leverages diffusion models to generate synthetic data that can be used to train segmentation models that transfer zero-shot to new classes (Li et al., 2023b). While effective, these methods are training-based and thus do not provide as much insight into the representations of existing text-to-image generation models, which is the key motivation behind ConceptAttention.
@@
Another line of work attempts perform unsupervised segmentation without any class or text conditioning by performing clustering of the embeddings of models (Cho et al., 2021; Hamilton et al., 2022; Tian et al., 2024). Despite not producing class predictions, these models are often evaluated on semantic segmentation datasets by using approaches like Hungarian matching (Kuhn, 1955) to pair unlabeled segmentation predictions with the best matching ones in a multi-class semantic segmentation dataset. In contrast, ConceptAttention enables text conditioning so we do not compare to this family of methods. We also don’t compare to models like SAM (Kirillov et al., 2023; Ravi et al., 2024) as it is trained on a large scale dataset.
@@ 4.1 Using ConceptAttention
$k_{c}$ $=[K_{p}c_{1},\dots K_{p}c_{k}]$ (4)
$q_{c}$ $=[Q_{p}c_{1},\dots Q_{p}c_{k}]$ (5)
$v_{c}$ $=[V_{p}c_{1},\dots V_{p}c_{k}]$ (6)
each in $\mathbb{R}^{r\times d}$.
@@ A Concept Residual Stream
The above operations create a residual stream of concept embeddings distinct from the image and patch embeddings. Following the pretrained transformer’s design, after the MMAttn we apply another projection matrix $P$ and MLP, adding the result residually to $c^{L}$. We apply an adaptive layernorm at the end of attention operation which outputs several values: a scale $\gamma$, shift $\beta$, and some gating values $\alpha_{1}$ and $\alpha_{2}$. The residual stream is then updated as
@@ 4.2 Limitations of Raw Cross Attention Maps
ImageNet-Segmentation PascalVOC (Single Class)
@@
LRP (Binder et al., 2016) CLIP ViT 51.09 32.89 55.68 48.77 31.44 52.89
DINO Attention (Caron et al., 2021) DINO ViT 81.97 69.44 86.12 80.71 64.33 88.90
DAAM (Tang et al., 2022) SDXL UNet 78.47 64.56 88.79 72.76 55.95 88.34
DAAM (Tang et al., 2022) SD2 UNet 64.52 47.62 78.01 64.28 45.01 83.04
Flux Cross Attention Flux DiT 74.92 59.90 87.23 80.37 54.77 89.08
ConceptAttention Flux DiT 83.07 71.04 90.45 87.85 76.45 90.19
@@
[…]
TextSpan 73.84 38.10
DAAM 62.89 10.97
Flux Cross Attention 79.52 27.04
ConceptAttention 86.99 51.39
5 Experiments
5.1 Zero-shot Image Segmentation
Datasets
Key Baseline Methods
We compare our approach to a variety of zero-shot interpretability methods which leverage several different multi-modal foundation models. We compare to numerous interpretability methods compatible with CLIP: Layerwise Relevance Propagation (LRP) (Binder et al., 2016), LRP on just the final-layer of a ViT (Partial-LRP), Attention Rollout (Rollout) (Abnar & Zuidema, 2020), Raw Vision Transformer Attention (ViT Attention) (Dosovitskiy et al., 2021), GradCAM (Selvaraju et al., 2019), TextSpan (Gandelsman et al., 2024), and the Transformer Attribution method from (Chefer et al., 2021) (TransInterp). We also compare to a UNet-based interpretability method that aggregates information from UNet cross attention layers called DAAM (Tang et al., 2022) on both Stable Diffusion XL (Podell et al., 2023) and Stable Diffusion 2. Finally, we compare to the raw cross attention maps produced by Flux DiT models.
Implementation Details
For all of our experiments we use the Flux DiT architecture implemented in PyTorch (Paszke et al., 2019). In particular, we use the distilled Flux-Schnell model. We encode real images with the DiT by first mapping them to the VAE latent space and then adding varying degrees of Gaussian noise before passing them through the Flux DiT. We then cache all of the concept output $o_{c}$ and image output vectors $o_{x}$ from each MMAttn layer. At the end of generation we then construct our concept saliency maps for each layer and average them over all layers of interest. In our experiments we leverage the activations from the last 10 of the 18 MMAttn layers.
@@ 5.1 Zero-shot Image Segmentation
Each method produces a set of scalar raw scores for each image patch which we then threshold based on the mean value to produce a binary segmentation prediction. We compare each method using standard segmentation evaluation metrics, namely: mean Intersection over Union (mIoU), patch/pixelwise accuracy (Acc), and mean Average Precision (mAP). Accuracy alone is an insufficient metric as our dataset is highly imbalanced, mIoU is significantly better, and mAP captures threshold agnostic segmentation capability. We found that ConceptAttention significantly out performs all of the baselines we tested across all three metrics (See Table 1). This is true for diffusion, CLIP, and DINO based interpretability methods.
@@
We also show qualitative results comparing the segmentation performance to each baseline in Figure 5. We also provide more qualitative results in Appendix B. It is worth noting that the qualitative segmentation results highlight (a) the ambiguity of zero-shot image segmentation, and (b) the limitations of human data annotation. For example, Figure 5 shows that our method does not segment the part of the dog between the ears and it’s body, while the data annotation does.
@@
We also wanted to evaluate the capabilities of our method at differentiating between multiple classes in an image. However, only a subset of methods produce distinct saliency maps for open ended classes. For this experiment we compare to DAAM using a SDXL backbone, TextSpan using a CLIP backbone, and the raw cross attentions of Flux. Instead of binarizing the image to produce segmentations, for each patch we predict the concept with the highest score. We used pixewlise accuracy and mIoU as our evaluation metrics and found that our method significantly out performed the baselines (See Table 4). We also show qualitative results of our approach differentiating between multiple concepts in a single image in Figures 1, 3 and we show more results in Appendix B.
@@ 5.2 Ablation Studies
We hypothesized that deeper MMAttn layers in the DiT would have more refined representations that better transfer to segmentation. This was confirmed by our evaluation (see Figure 6). We pull features from each diffusion layer and evaluated the segmentation performance of these features on ImageNet Segmentation. We also compare the performance of combining all layers simultaneously, which we found performs better than any individual layer.
Figure 6: Later MMAttn layers encode richer features for zero-shot segmentation. We investigated the impact of using features from various MMAttn layers and found that deeper layers lead to better performance on segmentation metrics like pixelwise accuracy, mIoU, and mAP. We also found that combining the information from all layers further improves performance.
@@
We add Gaussian noise to encoded images before passing them to the DiTs, this conforms with the expectations of the models. Intuitively one might expect the later timesteps (less noise) to have much higher segmentation performance as less information about the original image is corrupted. However, we found that the middle diffusion timesteps best (See Figure 7). Throughout the rest of our experiments we use timestep 500 out of 1000 following this result.
Figure 7: Optimal segmentation performance requires some noise to be present in the image. We evaluated the performance of ConceptAttention by encoding samples from a variety of timesteps (determines the amount of noise). Interestingly, we found that the optimal amount of noise was not zero, but in the middle to later stages of the noise schedule.
@@
We compared the performance on the ImageNet Segmentation benchmark of performing (a) just cross attention from the image patches to the concept vectors, (b) just self attention, (c) no attention operations, and (d) both cross and self attention. Our results seen in Table 3 indicate that using a combination of both cross and self attention achieves the best performance.
We also investigated the impact of applying a pixelwise softmax operation over our predicted segmentation coefficients. We found that it slightly improves segmentation performance in the attention output space and significantly improves performance for the cross attention maps (see Table 2.

### v2

@@ Abstract
Do the rich representations of multi-modal diffusion transformers (DiTs) exhibit unique properties that enhance their interpretability? We introduce ConceptAttention, a novel method that leverages the expressive power of DiT attention layers to generate high-quality saliency maps that precisely locate textual concepts within images. Without requiring additional training, ConceptAttention repurposes the parameters of DiT attention layers to produce highly contextualized concept embeddings, contributing the major discovery that performing linear projections in the output space of DiT attention layers yields significantly sharper saliency maps compared to commonly used cross-attention maps. ConceptAttention even achieves state-of-the-art performance on zero-shot image segmentation benchmarks, outperforming 15 other zero-shot interpretability methods on the ImageNet-Segmentation dataset. ConceptAttention works for popular image models and even seamlessly generalizes to video generation. Our work contributes the first evidence that the representations of multi-modal DiTs are highly transferable to vision tasks like segmentation.
@@ 1 Introduction
Diffusion models have recently gained widespread popularity, emerging as the state-of-the-art approach for a variety of generative tasks, particularly text-to-image synthesis (Rombach et al., 2022). These models transform random noise into photorealistic images guided by textual descriptions, achieving unprecedented fidelity and detail. Despite the impressive generative capabilities of diffusion models, our understanding of their internal mechanisms remains limited. Diffusion models operate as black boxes, where the relationships between input prompts and generated outputs are visible, but the decision-making processes that connect them are hidden from human understanding. footnotetext: Code: alechelbling.com/ConceptAttention/
@@
  • The novel discovery that the output vectors of attention operations produce higher-quality saliency maps than cross attentions. ConceptAttention repurposes the parameters of DiT attention layers to produce rich textual embeddings corresponding to different concepts, something that is uniquely enabled by multi-modal DiT architectures. By performing linear projections between these concept embeddings and image patch representations in the attention output space we can produce high quality saliency maps.
  • ConceptAttention achieves state-of-the-art performance in zero-shot segmentation on benchmarks like ImageNet Segmentation and Pascal VOC across multiple DiT architectures. We achieve superior performance to a diverse set of zero-shot interpretability methods based on various foundation models like CLIP, DINO, and UNet-based diffusion models; this highlights the potential for the representations of DiTs to transfer to important downstream vision tasks like segmentation. We replicate our results quantitatively on both Flux and Stable Diffusion 3.5 Turbo.
  • ConceptAttention works with a video generation DiT model. Additionally, we demonstrate qualitatively that ConceptAttention seamlessly generalizes to the CogVideoX (Yang et al., 2025) video generation MMDiT model, producing higher-quality saliency maps than native cross attention maps.
@@ 2 Related Work
A variety of works train models from scratch for the task of image segmentation (Amit et al., 2022; Karazija et al., 2024) or attempt to fine-tune pretrained models (Baranchuk et al., 2022). Another line of work leverages diffusion models to generate synthetic data that can be used to train segmentation models that transfer zero-shot to new classes (Li et al., 2023b). While effective, these methods are training-based and thus do not provide as much insight into the representations of existing text-to-image generation models, which is the key motivation behind ConceptAttention.
@@
Another line of work performs unsupervised segmentation without any class or text conditioning by performing clustering of the embeddings of models (Cho et al., 2021; Hamilton et al., 2022; Tian et al., 2024). Despite not producing class predictions, these models are often evaluated on semantic segmentation datasets by using approaches like Hungarian matching (Kuhn, 1955) to pair unlabeled segmentation predictions with the best matching ones in a multi-class semantic segmentation dataset. In contrast, ConceptAttention enables text conditioning so we do not compare to this family of methods. We also don’t compare to models like SAM (Kirillov et al., 2023; Ravi et al., 2024) as it is trained on a large scale dataset.
@@ 4.1 Using ConceptAttention
ImageNet-Segmentation PascalVOC (Single Class)
[…]
LRP (Binder et al., 2016) CLIP ViT 51.09 32.89 55.68 48.77 31.44 52.89
CLIPasRNN (Sun et al., 2024) CLIP ViT 74.05 58.80 84.80 61.76 41.48 76.57
OVAM (Marcos-Manchón et al., 2024) SDXL UNet 79.41 65.02 88.12 73.50 58.12 87.91
DINO SA (Caron et al., 2021) DINO ViT 81.97 69.44 86.12 80.71 64.33 88.90
DINOv2 SA (Oquab et al., 2024) DINOv2 ViT 77.39 63.12 84.19 79.65 57.61 87.26
DINOv2 Reg SA (Darcet et al., 2024) DINOv2 Reg 72.04 56.31 80.83 77.16 56.60 86.35
iBOT SA (Zhou et al., 2022) iBOT ViT 76.34 61.73 82.04 74.96 55.80 85.26
DAAM (Tang et al., 2022) SDXL UNet 78.47 64.56 88.79 72.76 55.95 88.34
DAAM (Tang et al., 2022) SD2 UNet 64.52 47.62 78.01 64.28 45.01 83.04
Cross Attention Flux DiT 74.92 59.90 87.23 80.37 54.77 89.08
Cross Attention SD3.5 DiT 77.80 63.67 83.50 80.22 61.46 86.97
ConceptAttention SD3.5 DiT 81.92 67.47 90.79 83.90 69.93 90.02
ConceptAttention Flux DiT 83.07 71.04 90.45 87.85 76.45 90.19
@@
$k_{c}=[K_{p}c_{1},\dots],q_{c}=[Q_{p}c_{1},\dots],v_{c}$ $=[V_{p}c_{1},\dots]\in\mathbb{R}^{r\times d}.$ (4)
@@ A Concept Residual Stream
The above operations create a residual stream of concept embeddings distinct from the image and patch embeddings. Following the pretrained transformer’s design, after the MMAttn we apply another projection matrix $P$ and MLP, adding the result residually to $c^{L}$. We apply an adaptive layernorm at the end of the attention operation which outputs several values: a scale $\gamma$, shift $\beta$, and some gating values $\alpha_{1}$ and $\alpha_{2}$. The residual stream is then updated as
@@ 4.2 Limitations of Raw Cross Attention Maps
Figure 6: ConceptAttention is capable of generating high quality saliency maps with Stable Diffusion 3.5 Turbo. Furthermore, the top example highlights a potential failure case of ConceptAttention. The concepts “sky”, “mountain”, and “sun” all semantically overlap, resulting in unclear object boundaries.
@@
TextSpan 73.84 38.10
DAAM 62.89 10.97
Flux Cross Attention 79.52 27.04
ConceptAttention 86.99 51.39
5 Experiments
5.1 Implementation Details
Flux DiT
For most of our experiments we use the Flux DiT architecture implemented in PyTorch (Paszke et al., 2019). In particular, we use the distilled Flux-Schnell model. We encode real images with the DiT by first mapping them to the VAE latent space and then adding varying degrees of Gaussian noise before passing them through the Flux DiT. We then cache all of the concept output $o_{c}$ and image output vectors $o_{x}$ from each MMAttn layer. At the end of generation we then construct our concept saliency maps for each layer and average them over all layers of interest. In our experiments we leverage the activations from the last 10 of the 18 MMAttn layers.
Stable Diffusion 3.5 Turbo
We found that our approach replicated on the Stable Diffusion 3.5 Turbo (Esser et al., 2024) DiT architecture (Figure 6).
CogVideo X
ConceptAttention generalizes to the CogVideoX (Yang et al., 2025) multi-modal DiT video generation model. The only change we make is additionally averaging over the added frame dimension.
5.2 Zero-shot Image Segmentation
Figure 7: (Left) Later MMAttn layers encode richer features for zero-shot segmentation. We investigated the impact of using features from various MMAttn layers and found that deeper layers lead to better performance on segmentation metrics like pixelwise accuracy, mIoU, and mAP. We also found that combining the information from all layers further improves performance. (Right) Optimal segmentation performance requires some noise to be present in the image. We evaluated the performance of ConceptAttention by encoding samples from a variety of timesteps (determines the amount of noise). Interestingly, we found that the optimal amount of noise was not zero, but in the middle to later stages of the noise schedule.
@@ 5.2 Zero-shot Image Segmentation
We compare our approach to a variety of zero-shot interpretability methods which leverage several different multi-modal foundation models. We compare to numerous interpretability methods compatible with CLIP: Layerwise Relevance Propagation (LRP) (Binder et al., 2016), LRP on just the final-layer of a ViT (Partial-LRP), Attention Rollout (Rollout) (Abnar & Zuidema, 2020), Raw Vision Transformer Attention (ViT Attention) (Dosovitskiy et al., 2021), GradCAM (Selvaraju et al., 2019), TextSpan (Gandelsman et al., 2024), CLIP as RNN (Sun et al., 2024), and the Transformer Attribution method from (Chefer et al., 2021) (TransInterp). We also compare to UNet-based interpretability methods that aggregates information from UNet cross attention layers called DAAM (Tang et al., 2022) on both SDXL (Podell et al., 2023) and SD2, and OVAM (Li et al., 2023b) with SDXL. We compare to the self-attention maps of various DINO models: DINOv1 (Caron et al., 2021), DINOv2 (Oquab et al., 2024), and DINOv2 with registers (Darcet et al., 2024). Finally, we compare to the raw cross attention maps produced by Flux and Stable Diffusion 3.5 Turbo.
@@
Each method produces a set of scalar raw scores for each image patch which we then threshold based on the mean value to produce a binary segmentation prediction. We compare each method using standard segmentation evaluation metrics, namely: mean Intersection over Union (mIoU), patch/pixelwise accuracy (Acc), and mean Average Precision (mAP). Accuracy alone is an insufficient metric as our dataset is highly imbalanced, mIoU is significantly better, and mAP captures threshold agnostic segmentation capability. We found that ConceptAttention significantly out performs all of the baselines we tested across all three metrics (Table 1). This is true for diffusion, CLIP, and DINO based interpretability methods.
@@
We show qualitative results comparing the segmentation performance to each baseline in Figure 4 and more qualitative results in Appendix B. It is worth noting that the qualitative segmentation results highlight (a) the ambiguity of zero-shot image segmentation, and (b) the limitations of human data annotation. For example, Figure 4 shows our method does not segment the part of the dog between the ears and it’s body, while the ground truth does.
@@
Figure 8: ConceptAttention generalizes seamlessly to video generation MMDiT models like CogVideoX. We apply ConceptAttention to a CogVideoX (Yang et al., 2025) video generation model. We take several frames from the video and compare the saliency maps generated by ConceptAttention to the model’s internal cross attention maps.
We also wanted to evaluate the capabilities of our method at differentiating between multiple classes in an image. However, only a subset of methods produce distinct saliency maps for open ended classes. For this experiment we compare to DAAM using a SDXL backbone, TextSpan using a CLIP backbone, and the raw cross attentions of Flux. Instead of binarizing the image to produce segmentations, for each patch we predict the concept with the highest score. We used pixelwise accuracy and mIoU as our evaluation metrics and found that our method significantly outperformed the baselines (Table 2). We also show qualitative results of our approach differentiating between multiple concepts in a single image in Figures 1, 3 and we show more results in Appendix B.
@@ 5.3 Ablation Studies
We hypothesized that deeper MMAttn layers in the DiT would have more refined representations that better transfer to segmentation. This was confirmed by our evaluation (Figure 7). We pull features from each diffusion layer and evaluated the segmentation performance on ImageNet-Segmentation. We also compare the performance of combining all layers simultaneously, which we found performs better than any individual layer.
@@
We add Gaussian noise to encoded images before passing them to the DiTs, this conforms with the expectations of the models. Intuitively one might expect the later timesteps (less noise) to have much higher segmentation performance as less information about the original image is corrupted. However, we found that the middle diffusion timesteps best (Figure 7). Throughout the rest of our experiments we use timestep 500 out of 1000 following this result.
@@
We compared the performance on the ImageNet Segmentation benchmark of performing (a) just cross attention from the image patches to the concept vectors, (b) just self attention, (c) no attention operations, and (d) both cross and self attention. Our results seen in Table 4 indicate that using a combination of both cross and self attention achieves the best performance. We also investigated the impact of applying a pixelwise softmax operation over our predicted segmentation coefficients. We found that it slightly improves segmentation performance in the attention output space and significantly improves performance for the cross attention maps (Table 3)
5.4 Video Model Results
We include qualitative results demonstrating the efficacy of ConceptAttention on the CogVideoX video generation multi-modal DiT model (Figure 8). Also see Figures 16 and 17 in the Appendix.
5.5 Limitations
The primary limitation of ConceptAttention is that it struggles to differentiate between very similar textual concepts. For example, for a photo with a sky with the sun in it, the model does not necessarily know where the boundary of the sun resides, instead capturing a halo around the sun (Figure 6). Additionally, when no relevant concept is present, ConceptAttention will select the most similar one even if it is incorrect (Figure 15 in the Appendix).
@@ Appendix B More Qualitative Results
Figure 15: The behavior of ConceptAttention when multiple relevant concepts are present and when no relevant one is. When multiple similar concepts are given, like “car” and “bike”, the most similar one will be chosen. However, when no relevant concept is presented, ConceptAttention will fall back on the most relevant one, in this case “car“ for the bike patches.
Appendix C Concept Attention on Video Generation Models
Figure 16: ConceptAttention generalizes seamlessly to video generation MMDiT models like CogVideoX. We apply ConceptAttention to a CogVideoX (Yang et al., 2025) video generation model. We take several frames from the video and compare the saliency maps generated by ConceptAttention to the model’s internal cross attention maps.
Figure 17: ConceptAttention generalizes seamlessly to video generation MMDiT models like CogVideoX. We apply ConceptAttention to a CogVideoX (Yang et al., 2025) video generation model. We take several frames from the video and compare the saliency maps generated by ConceptAttention to the model’s internal cross attention maps.